### v2

# Closed-Form Evaluation of $\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ via Infinite Products

Abstract
We establish closed-form expressions for the infinite series $\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ for all integers $k\geq 2$ by connecting these sums to infinite product formulas involving the gamma function. Our approach uses logarithmic manipulations, the Fubini–Tonelli theorem, and Frullani’s integral theorem. As applications, we derive a structural identity relating the Riemann zeta function $\zeta(k)$ to these sums, establish a new series representation for the Euler–Mascheroni constant $\gamma$, and show that this representation admits an exponentially convergent reformulation via zeta values. We further prove that the sum $h(k)=\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ is a strictly decreasing, strictly convex function of $k$ and establish explicit two-sided bounds and asymptotic expansions. The decimal expansions of the closed-form values and several auxiliary sequences arising from these identities are cataloged in the OEIS.
1 Introduction
The evaluation of infinite series involving inverse hyperbolic functions has a rich history in analysis and number theory. While series of the form $\sum_{n=1}^{\infty}n^{-s}$ are thoroughly understood through the theory of the Riemann zeta function, series involving $\operatorname{arctanh}(n^{-k})$ have received comparatively less systematic treatment.
In this paper, we establish explicit closed-form expressions for
$h(k):=\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right),\qquad k\geq 2.$ (1)
Our approach connects these sums to certain infinite products whose exact values were determined by Cantrell [2] in terms of the gamma function and elementary trigonometric functions; these product identities rest ultimately on the Weierstrass factorization theorem for $1/\Gamma(z)$ and the reflection and multiplication formulas for the gamma function (see, e.g., Prudnikov, Brychkov, and Marichev [4], Borwein, Bailey, and Girgensohn [1], and the NIST Digital Library [3]).
The sequence of values $h(2),h(3),h(4),\ldots$ begins
$0.650923\ldots,\quad 0.202733\ldots,\quad 0.082405\ldots,\quad 0.036938\ldots,\quad 0.017344\ldots\,$,
and the decimal expansion of $h(2)=\tfrac{1}{2}\ln(\sinh(\pi)/\pi)$ appears in the OEIS as A372862.
1.1 Main results
Our principal result is the following closed-form evaluation.
Theorem 1.
For all integers $k\geq 2$,
$\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right)=\frac{1}{2}\ln\!\left(\frac{f(k)}{2\,g(k)}\right),$ (2)
where
$f(k):=\prod_{n=1}^{\infty}\!\left(1+\frac{1}{n^{k}}\right)\quad\text{and}\quad g(k):=\prod_{n=2}^{\infty}\!\left(1-\frac{1}{n^{k}}\right)$ (3)
are given by explicit formulas in terms of the gamma function (for odd $k$) or trigonometric functions (for even $k$); see Theorem 5.
The sum $h(k)$ admits the equivalent zeta-series representation
$h(k)=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1},$ (4)
which follows from the Taylor expansion of $\operatorname{arctanh}$ and the Fubini–Tonelli theorem (see Remark 9).
As consequences, we obtain a structural identity for the Riemann zeta function and a new representation of the Euler–Mascheroni constant.
Corollary 2.
For $k\geq 2$,
$\zeta(k)=1+h(k)-C(k),$ (5)
where
$C(k):=\sum_{n=1}^{\infty}\frac{\zeta\bigl(k(2n+1)\bigr)-1}{2n+1}$ (6)
is a rapidly convergent correction satisfying $C(k)\to 0$ as $k\to\infty$.
Corollary 3.
The Euler–Mascheroni constant $\gamma$ satisfies
$\gamma=1-\frac{1}{2}\ln 2-\sum_{n=2}^{\infty}\!\left(\operatorname{arctanh}\!\left(\frac{1}{n}\right)-\frac{1}{n}\right).$ (7)
1.2 Related work
Talla Waffo [5] recently investigated the Mellin transforms of $1/\!\operatorname{arctanh}(x)$ and related functions on $(0,1)$, establishing connections to derivatives of the Dirichlet eta and beta functions at negative integers. While the continuous integral approach of [5] and our discrete summation approach are technically independent, both highlight the rich analytic structure underlying inverse hyperbolic functions.
1.3 Structure of the paper
Section 2 recalls the infinite product formulas and establishes preliminary results. Section 3 contains the proof of Theorem 1. Section 4 derives the zeta function identity and studies the correction function $C(k)$. Section 5 establishes the representation for $\gamma$ using Frullani’s theorem. Section 6 develops the analytic properties of $h(k)$, including monotonicity, convexity, asymptotic behavior, and two-sided bounds. Section 7 presents an exponentially convergent reformulation of the $\gamma$ representation and establishes tail bounds for the direct computation of $h(k)$. Section 8 collects numerical examples.
2 Infinite product formulas
We begin by recalling the product formulas that form the foundation of our approach.
Definition 4.
For integers $k\geq 2$, define
$f(k):=\prod_{n=1}^{\infty}\!\left(1+\frac{1}{n^{k}}\right)\quad\text{and}\quad g(k):=\prod_{n=2}^{\infty}\!\left(1-\frac{1}{n^{k}}\right).$ (8)
These products converge for $k\geq 2$, since taking logarithms yields series that converge by comparison with $\sum n^{-k}$.
Theorem 5 (Cantrell [2]; cf. [4, 1, 3]).
For integers $k\geq 2$, the products $f(k)$ and $g(k)$ have the following closed-form evaluations.
  • (a)
    For odd $k$:
    $f(k)$ $=\frac{1}{\prod_{j=1}^{k-1}\Gamma\bigl[(-1)^{j}(1+1/k)\bigr]},$ (9)
    $g(k)$ $=\frac{1}{k\prod_{j=1}^{k-1}\Gamma\bigl[(-1)^{1+j}(1+1/k)\bigr]}.$ (10)
  • (b)
    For even $k$:
    $f(k)$ $=\frac{\prod_{j=1}^{k/2}\sin\!\bigl[\pi(-1)^{(2j-1)/k}\bigr]}{(\pi i)^{k/2}},$ (11)
    $g(k)$ $=\frac{\prod_{j=1}^{(k/2)-1}\sin\!\bigl[\pi(-1)^{2j/k}\bigr]}{k(\pi i)^{(k/2)-1}}.$ (12)
The proof relies on the Weierstrass product representation $1/\Gamma(z)=ze^{\gamma z}\prod_{n=1}^{\infty}(1+z/n)e^{-z/n}$, the Euler reflection formula $\Gamma(z)\Gamma(1-z)=\pi/\sin(\pi z)$, and the Gauss multiplication formula for $\Gamma$. These product evaluations were communicated by Cantrell to MathWorld [2] and follow from the general principle that convergent infinite products whose factors are rational functions of the index can always be expressed in terms of finite products of gamma values (Borwein, Bailey, and Girgensohn [1]; Prudnikov, Brychkov, and Marichev [4]).
Example 6.
For $k=2$, we have $f(2)=\sinh(\pi)/\pi$ and $g(2)=1/2$, giving
$h(2)=\frac{1}{2}\ln\!\left(\frac{\sinh(\pi)}{\pi}\right)\approx 0.650923.$
The decimal expansion of this constant appears in the OEIS as A372862.
To connect these products to our desired sums, we require the following lemmas relating logarithms of products to series involving the Riemann zeta function.
Lemma 7.
For $k\geq 2$,
$\ln f(k)=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}\zeta(kn)}{n}.$ (13)
Proof.
Taking logarithms in the definition of $f(k)$ and applying the Taylor series $\ln(1+x)=\sum_{m=1}^{\infty}(-1)^{m+1}x^{m}/m$ with $x=n^{-k}$ yields a double series. Since $k\geq 2$, absolute convergence holds, and the Fubini–Tonelli theorem permits interchange of summation:
$\ln f(k)=\sum_{n=1}^{\infty}\sum_{m=1}^{\infty}\frac{(-1)^{m+1}}{m\,n^{km}}=\sum_{m=1}^{\infty}\frac{(-1)^{m+1}}{m}\sum_{n=1}^{\infty}\frac{1}{n^{km}}=\sum_{m=1}^{\infty}\frac{(-1)^{m+1}\zeta(km)}{m}.\qed$
Lemma 8.
For $k\geq 2$,
$\ln g(k)=-\sum_{n=1}^{\infty}\frac{\zeta(kn)-1}{n}.$ (14)
Proof.
The proof is analogous, using $\ln(1-x)=-\sum_{m=1}^{\infty}x^{m}/m$ and noting that $\sum_{n=2}^{\infty}n^{-km}=\zeta(km)-1$. ∎
3 Proof of the main theorem
Proof of Theorem 1.
Subtracting the identity in Lemma 8 from that in Lemma 7 gives
$\ln\!\left(\frac{f(k)}{g(k)}\right)=\sum_{n=1}^{\infty}\frac{\bigl[(-1)^{n+1}+1\bigr]\zeta(kn)-1}{n}.$
The coefficient $(-1)^{n+1}+1$ equals $2$ when $n$ is odd and 0 when $n$ is even. Therefore
$\ln\!\left(\frac{f(k)}{g(k)}\right)=\sum_{n=1}^{\infty}\frac{2\bigl[\zeta\bigl(k(2n-1)\bigr)-1\bigr]}{2n-1}+\sum_{n=1}^{\infty}\frac{1}{2n-1}.$ (15)
Because the second sum is divergent, we regroup in adjacent odd-even blocks without reordering terms to keep all sums convergent:
$\ln\!\left(\frac{f(k)}{g(k)}\right)=2\zeta(k)-1-\sum_{n=1}^{\infty}\left[\frac{1}{2n}+\frac{1-2\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}\right].$
Now, for the sum portion,
$\sum_{n=1}^{\infty}\left[\frac{1}{2n}+\frac{1-2\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}\right]$ $=\sum_{n=1}^{\infty}\left[\frac{2n+1}{(2n)(2n+1)}+\frac{2n\!\left(1-2\zeta\!\bigl(k(2n+1)\bigr)\right)}{(2n)(2n+1)}\right]$
$=\sum_{n=1}^{\infty}\frac{1}{2n(2n+1)}+2\sum_{n=1}^{\infty}\frac{1-\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}.$
Also,
$\sum_{n=1}^{\infty}\frac{1}{2n(2n+1)}=\sum_{n=1}^{\infty}\left(\frac{1}{2n}-\frac{1}{2n+1}\right)=1-\ln 2.$
Hence
$\ln\!\left(\frac{f(k)}{g(k)}\right)=2\zeta(k)-2+\ln 2-2\sum_{n=1}^{\infty}\frac{1-\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}.$ (16)
For $s>1$, we have $1-\zeta(s)=-\sum_{m=2}^{\infty}m^{-s}$. By the Fubini–Tonelli theorem,
$\sum_{n=1}^{\infty}\frac{1-\zeta\bigl(k(2n+1)\bigr)}{2n+1}=-\sum_{m=2}^{\infty}\sum_{n=1}^{\infty}\frac{1}{(2n+1)\,m^{k(2n+1)}}.$ (17)
The key observation is that for $|x|<1$, $\operatorname{arctanh}(x)=\sum_{n=0}^{\infty}x^{2n+1}/(2n+1)$, so
$\sum_{n=1}^{\infty}\frac{(m^{-k})^{2n+1}}{2n+1}=\operatorname{arctanh}(m^{-k})-m^{-k}.$
Substituting into (17) and then into (16) gives
$\ln\!\left(\frac{f(k)}{g(k)}\right)$ $=2\zeta(k)-2+\ln 2+2\sum_{m=2}^{\infty}\bigl[\operatorname{arctanh}(m^{-k})-m^{-k}\bigr]$
$=2\zeta(k)-2+\ln 2+2\sum_{m=2}^{\infty}\operatorname{arctanh}(m^{-k})-2\bigl[\zeta(k)-1\bigr]$
$=\ln 2+2\sum_{m=2}^{\infty}\operatorname{arctanh}(m^{-k}).$
Therefore
$\frac{1}{2}\ln\!\left(\frac{f(k)}{2\,g(k)}\right)=\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right),$
which yields the result. ∎
Remark 9.
For every integer $k\geq 2$, the sum $h(k)$ admits the equivalent representations
$h(k)$ $=\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right)$ (18)
$=\frac{1}{2}\ln\!\left(\frac{f(k)}{2\,g(k)}\right)$ (19)
$=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}.$ (20)
Here (19) is Theorem 1, while (20) follows from the Taylor expansion $\operatorname{arctanh}(x)=\sum_{m=0}^{\infty}x^{2m+1}/(2m+1)$ applied with $x=n^{-k}$ and exchange of summation (justified by absolute convergence for $k\geq 2$).
4 Connection to the Riemann zeta function
Proof of Corollary 2.
From the proof of Theorem 1, we obtained
$\frac{1}{2}\ln\!\left(\frac{f(k)}{g(k)}\right)=\zeta(k)-1+\sum_{n=1}^{\infty}\frac{\zeta\bigl(k(2n+1)\bigr)-1}{2n+1}+1-\ln 2+\frac{1}{2}.$
More directly, combining (20) with the definition of $C(k)$ gives
$h(k)=\bigl[\zeta(k)-1\bigr]+C(k),$
so $\zeta(k)=1+h(k)-C(k)$. ∎
4.1 Properties of the correction function
Proposition 10.
The correction function $C(k)$ satisfies the following.
  • (a)
    $C(k)>0$ for all $k\geq 2$.
  • (b)
    $C(k)\to 0$ as $k\to\infty$.
  • (c)
    $C(k)$ is a strictly decreasing function of $k$.
Proof.
Part (a) holds because $\zeta(s)>1$ for all $s>1$. Part (b) follows from $\zeta\bigl(k(2n+1)\bigr)\to 1$ exponentially fast as $k\to\infty$ and dominated convergence. For part (c), if $k_{1}<k_{2}$ then $k_{1}(2n+1)<k_{2}(2n+1)$, so $\zeta\bigl(k_{1}(2n+1)\bigr)>\zeta\bigl(k_{2}(2n+1)\bigr)$, whence $C(k_{1})>C(k_{2})$. ∎
Theorem 11.
For integers $k\geq 2$, the error in the approximation $\zeta(k)\approx 1+h(k)$ satisfies
$\bigl|\zeta(k)-1-h(k)\bigr|=C(k)\leq\frac{1}{3}\!\left(1+\frac{2}{3k-1}\right)\frac{2^{-3k}}{1-2^{-2k}}.$ (21)
In particular, $C(k)=O(2^{-3k})$ as $k\to\infty$.
Proof.
For $s>1$, we have $\zeta(s)-1=\sum_{m=2}^{\infty}m^{-s}\leq 2^{-s}+\int_{2}^{\infty}x^{-s}\,dx=2^{-s}(1+2/(s-1))$. Writing $s_{n}=k(2n+1)$, we have $s_{n}\geq 3k$ and $2n+1\geq 3$ for $n\geq 1$, so
$C(k)\leq\frac{1}{3}\!\left(1+\frac{2}{3k-1}\right)\sum_{n=1}^{\infty}2^{-k(2n+1)}=\frac{1}{3}\!\left(1+\frac{2}{3k-1}\right)\frac{2^{-3k}}{1-2^{-2k}}.\qed$
Remark 12.
The identity $\zeta(k)=1+h(k)-C(k)$ is of structural rather than computational interest: it decomposes $\zeta(k)$ into a dominant $\operatorname{arctanh}$-sum plus a correction that decays as $O(2^{-3k})$. For $k=2$, the approximation $\zeta(2)\approx 1+h(2)\approx 1.65092$ already achieves a relative error less than $0.001\%$ from the exact value $\pi^{2}/6\approx 1.64493$.
5 Representation for the Euler–Mascheroni constant
We now derive a new series representation for the Euler–Mascheroni constant using properties of the correction function and Frullani’s integral theorem.
5.1 Derivative of the correction function
Lemma 13.
The derivative of $C(k)$ is given by
$C^{\prime}(k)=-\sum_{m=2}^{\infty}\frac{\ln m}{m^{3k}-m^{k}}.$ (22)
Proof.
Differentiating the definition (6) term-by-term and using $\zeta^{\prime}(s)=-\sum_{m=2}^{\infty}(\ln m)/m^{s}$, we obtain $C^{\prime}(k)=\sum_{n=1}^{\infty}\zeta^{\prime}\bigl(k(2n+1)\bigr)$. Exchanging the order of summation (justified by absolute convergence for $k\geq 1$) and evaluating the resulting geometric series yields (22). ∎
5.2 Integration and application of Frullani’s theorem
Lemma 14.
For $m\geq 2$ and $k\geq 1$,
$\int_{1}^{k}\frac{\ln m}{m^{3t}-m^{t}}\,dt=m^{-k}-\operatorname{arctanh}(m^{-k})-m^{-1}+\operatorname{arctanh}(m^{-1}).$
Proof.
We use the partial fraction decomposition $1/(m^{3t}-m^{t})=\frac{1}{2}\bigl[1/(m^{t}-1)-1/(m^{t}+1)\bigr]/m^{t}$ and the substitution $u=m^{-t}$ in each integral. Computing the resulting logarithms and expressing them in terms of $\operatorname{arctanh}$ gives the stated formula. ∎
Proposition 15.
The correction function at $k=1$ satisfies
$C(1)=1-\gamma-\tfrac{1}{2}\ln 2.$ (23)
Proof.
From the definition, $C(1)=\sum_{n=1}^{\infty}[\zeta(2n+1)-1]/(2n+1)$. Using the integral representation $\zeta(2n+1)-1]/(2n+1)=\int_{0}^{\infty}x^{2n}/[(2n+1)!\,e^{x}(e^{x}-1)]\,dx$, we write
$C(1)=\int_{0}^{\infty}\frac{\sinh(x)-x}{x\,e^{x}(e^{x}-1)}\,dx.$
Decomposing the integrand and applying Frullani’s theorem to the integral $\int_{0}^{\infty}(e^{-2x}-e^{-x})/x\,dx=-\frac{1}{2}\ln 2$ (with $f(x)=e^{-x}$, $a=2$, $b=1$), and using the standard representation $\gamma=\int_{0}^{\infty}[1/(e^{x}-1)-1/(xe^{x})]\,dx$, we obtain $C(1)=1-\gamma-\frac{1}{2}\ln 2$. ∎
5.3 Proof of the gamma formula
Proof of Corollary 3.
Integrating $C^{\prime}(k)$ from 11 to $k$ using Lemmas 13 and 14:
$C(k)-C(1)$ $=-\sum_{m=2}^{\infty}\int_{1}^{k}\frac{\ln m}{m^{3t}-m^{t}}\,dt$
$=\sum_{m=2}^{\infty}\bigl[\operatorname{arctanh}(m^{-k})-m^{-k}-\operatorname{arctanh}(m^{-1})+m^{-1}\bigr].$
Rearranging and using the definitions of $h(k)$ and $\zeta(k)$:
$C(k)=h(k)-\zeta(k)+1-\sum_{m=2}^{\infty}\bigl[\operatorname{arctanh}(m^{-1})-m^{-1}\bigr]+C(1).$
Substituting the identity $\zeta(k)=1+h(k)-C(k)$ from Corollary 2, all $k$-dependent terms cancel, leaving
$\sum_{m=2}^{\infty}\!\left(\frac{1}{m}-\operatorname{arctanh}\!\left(\frac{1}{m}\right)\right)=C(1).$
By Proposition 15, $C(1)=1-\gamma-\frac{1}{2}\ln 2$, and solving for $\gamma$ yields (7). ∎
Remark 16.
The series in (7) converges, but slowly: the general term $\operatorname{arctanh}(1/n)-1/n\sim 1/(3n^{3})$ as $n\to\infty$, so the tail after $N$ terms is $O(N^{-2})$. In Section 7, we show that expanding each $\operatorname{arctanh}$ term and resumming yields an equivalent formula involving $\zeta(2m+1)$ that converges geometrically, dramatically improving the rate from polynomial to exponential.
6 Analytic properties of $h(k)$
In this section, we study the behavior of $h(k)$ as a function of the continuous variable $k>1$. We establish monotonicity, convexity, an asymptotic expansion, and explicit two-sided bounds.
6.1 Monotonicity and convexity
Theorem 17.
For $k>1$, the function $h(k)=\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ is strictly decreasing and strictly convex. More precisely,
$h^{\prime}(k)$ $=-\sum_{n=2}^{\infty}\frac{(\ln n)\,n^{-k}}{1-n^{-2k}}<0,$ (24)
$h^{\prime\prime}(k)$ $=\sum_{n=2}^{\infty}\frac{(\ln n)^{2}\,n^{-k}(1+n^{-2k})}{(1-n^{-2k})^{2}}>0.$ (25)
Proof.
Set $\varphi_{n}(k):=\operatorname{arctanh}(n^{-k})$ for $n\geq 2$. With $t=n^{-k}$ we have $dt/dk=-(\ln n)\,n^{-k}$ and $d^{2}t/dk^{2}=(\ln n)^{2}\,n^{-k}$, so the chain rule gives
$\varphi_{n}^{\prime}(k)$ $=\frac{dt/dk}{1-t^{2}}=\frac{-(\ln n)\,n^{-k}}{1-n^{-2k}},$
$\varphi_{n}^{\prime\prime}(k)$ $=\frac{(d^{2}t/dk^{2})(1-t^{2})+2t(dt/dk)^{2}}{(1-t^{2})^{2}}=\frac{(\ln n)^{2}\,n^{-k}(1+n^{-2k})}{(1-n^{-2k})^{2}}.$
Every summand in (24) is strictly negative and every summand in (25) is strictly positive; the series converge uniformly on compact subsets of $(1,\infty)$ since $n^{-k}(\ln n)^{j}/(1-n^{-2k})^{j}$ is dominated by $C\,n^{-k+\varepsilon}$ for any $\varepsilon>0$. Term-by-term differentiation is therefore justified by the Weierstrass $M$-test. ∎
6.2 Asymptotic expansion
Theorem 18.
As $k\to\infty$,
$h(k)=\bigl(\zeta(k)-1\bigr)+\frac{\zeta(3k)-1}{3}+\frac{\zeta(5k)-1}{5}+\cdots=\sum_{m=0}^{M}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}+R_{M}(k),$ (26)
where the remainder satisfies $|R_{M}(k)|\leq\frac{4}{3}\cdot 2^{-(2M+3)k}$ for all $k\geq 2$. In particular,
$h(k)=\bigl(\zeta(k)-1\bigr)+O\bigl(2^{-3k}\bigr)=2^{-k}+3^{-k}+4^{-k}+O\bigl(5^{-k}+2^{-3k}\bigr).$ (27)
Proof.
The zeta-series representation (20) gives $h(k)=\sum_{m=0}^{\infty}\bigl[\zeta\bigl((2m+1)k\bigr)-1\bigr]/(2m+1)$. The remainder after $M$ terms is bounded using $\zeta(s)-1\leq 2^{-s}/(1-2^{-s})\leq 2^{1-s}$ for $s\geq 2$:
$|R_{M}(k)|\leq\sum_{m=M+1}^{\infty}\frac{2^{1-(2m+1)k}}{2m+1}<2\sum_{m=M+1}^{\infty}2^{-(2m+1)k}=\frac{2\cdot 2^{-(2M+3)k}}{1-2^{-2k}}\leq\frac{4}{3}\cdot 2^{-(2M+3)k},$
where the last inequality uses $1-2^{-2k}\geq 3/4$ for $k\geq 1$. Setting $M=0$ gives $h(k)=(\zeta(k)-1)+O(2^{-3k})$. ∎
6.3 Two-sided bounds
Theorem 19.
For $k\geq 2$,
$\zeta(k)-1<h(k)$ (28)
and
$h(k)<\zeta(k)-1+\frac{\zeta(3k)-1}{3(1-2^{-2k})}.$ (29)
More refined lower bounds are obtained by including additional terms from the zeta-series representation (20):
$h(k)>\zeta(k)-1+\frac{\zeta(3k)-1}{3}+\frac{\zeta(5k)-1}{5}.$ (30)
Proof.
For the lower bound (28), we use $\operatorname{arctanh}(x)>x$ for $0<x<1$, whence $h(k)>\sum_{n=2}^{\infty}n^{-k}=\zeta(k)-1$.
For (30), we simply note that all terms in the zeta-series (20) are positive, so any partial sum is a lower bound.
For the upper bound (29), the inequality $\operatorname{arctanh}(x)=x+\frac{x^{3}}{3}+\frac{x^{5}}{5}+\cdots<x+\frac{x^{3}}{3}\bigl(1+x^{2}+x^{4}+\cdots\bigr)=x+\frac{x^{3}}{3(1-x^{2})}$ holds for $0<x<1$. Setting $x=n^{-k}$ and summing over $n\geq 2$:
$h(k)<\bigl(\zeta(k)-1\bigr)+\frac{1}{3}\sum_{n=2}^{\infty}\frac{n^{-3k}}{1-n^{-2k}}.$
Since $n\geq 2$ implies $n^{-2k}\leq 2^{-2k}$, we have $1/(1-n^{-2k})\leq 1/(1-2^{-2k})$, giving (29). ∎
Example 20.
For $k=2$: the lower bound gives $h(2)>\zeta(2)-1=0.6449\ldots$, the refined lower bound gives $h(2)>0.6449+0.0060+0.0001=0.6510\ldots$, and the upper bound gives $h(2)<0.6449+0.0064=0.6514\ldots$. The exact value is $h(2)=0.6509\ldots$, which lies comfortably within these bounds.
6.4 Difference asymptotics
For integer sequences, the forward differences $\Delta h(k):=h(k)-h(k+1)$ carry structural information about the rate of decay.
Theorem 21.
For integers $k\geq 2$,
$\Delta h(k)=\sum_{n=2}^{\infty}n^{-k}\!\left(1-\frac{1}{n}\right)+O\bigl(2^{-3k}\bigr)=\bigl(\zeta(k)-\zeta(k\!+\!1)\bigr)+O\bigl(2^{-3k}\bigr).$ (31)
More explicitly, $\Delta h(k)$ admits the asymptotic expansion
$\Delta h(k)=2^{-(k+1)}+\frac{2}{3}\cdot 3^{-k}+\frac{3}{4}\cdot 4^{-k}+O\bigl(5^{-k}\bigr),$ (32)
and for any fixed $N\geq 2$,
$\Delta h(k)=\sum_{n=2}^{N}n^{-k}\!\left(1-\frac{1}{n}\right)+O\bigl((N\!+\!1)^{-k}+2^{-3k}\bigr).$ (33)
In particular, $\Delta h(k)\sim 2^{-(k+1)}$ as $k\to\infty$.
Proof.
From the zeta-series representation (20),
$\Delta h(k)=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-\zeta\bigl((2m+1)(k+1)\bigr)}{2m+1}.$
The $m=0$ term contributes $\zeta(k)-\zeta(k+1)=\sum_{n=2}^{\infty}n^{-k}(1-n^{-1})$. For $m\geq 1$, we bound
$\sum_{m=1}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-\zeta\bigl((2m+1)(k+1)\bigr)}{2m+1}\leq\sum_{m=1}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}=C(k)=O(2^{-3k}),$
using Theorem 11. The leading terms of $\sum_{n=2}^{\infty}n^{-k}(1-1/n)$ are $2^{-k}\cdot\frac{1}{2}+3^{-k}\cdot\frac{2}{3}+4^{-k}\cdot\frac{3}{4}+\cdots$, giving (32). ∎
Remark 22.
The sequence of differences $\Delta h(2),\Delta h(3),\Delta h(4),\ldots$ begins $0.4482,0.1203,0.0455,0.0196,0.0090,\ldots$ and the ratio $\Delta h(k)/2^{-(k+1)}$ converges to $1$ from above, consistent with the positive correction from $n=3,4,\ldots$ in (32). The decimal expansions of these values may be of independent interest for OEIS submission.
7 Accelerated computation
7.1 Exponentially convergent formula for $\gamma$
The series (7) for $\gamma$ converges only polynomially. We now show that the same identity, combined with the zeta-series representation, yields a formula with exponential convergence.
Theorem 23.
The Euler–Mascheroni constant satisfies
$\gamma=1-\frac{1}{2}\ln 2-\sum_{m=1}^{\infty}\frac{\zeta(2m+1)-1}{2m+1},$ (34)
where the series converges geometrically. The tail after $N$ terms satisfies
$0<\sum_{m=N+1}^{\infty}\frac{\zeta(2m+1)-1}{2m+1}<\frac{4}{3}\cdot 2^{-(2N+3)}.$ (35)
In particular, NN terms of (34) determine $\gamma$ to within $O(4^{-N})$, compared to $O(N^{-2})$ for the direct series (7).
Proof.
From (7),
$\gamma=1-\tfrac{1}{2}\ln 2-\sum_{n=2}^{\infty}\bigl[\operatorname{arctanh}(n^{-1})-n^{-1}\bigr].$
Expanding $\operatorname{arctanh}(n^{-1})-n^{-1}=\sum_{m=1}^{\infty}n^{-(2m+1)}/(2m+1)$ and exchanging the order of summation (justified by absolute convergence) gives
$\sum_{n=2}^{\infty}\bigl[\operatorname{arctanh}(n^{-1})-n^{-1}\bigr]=\sum_{m=1}^{\infty}\frac{1}{2m+1}\sum_{n=2}^{\infty}\frac{1}{n^{2m+1}}=\sum_{m=1}^{\infty}\frac{\zeta(2m+1)-1}{2m+1},$
which is precisely $C(1)$ from Proposition 15. This establishes (34).
For the tail bound, $\zeta(s)-1\leq 2^{1-s}$ for $s\geq 2$ (as in the proof of Theorem 18), so
$\sum_{m=N+1}^{\infty}\frac{\zeta(2m+1)-1}{2m+1}<2\sum_{m=N+1}^{\infty}2^{-(2m+1)}=\frac{2\cdot 2^{-(2N+3)}}{1-4^{-1}}=\frac{4}{3}\cdot 2^{-(2N+3)}.\qed$
Remark 24.
The acceleration is dramatic. Ten terms of (34) yield $\gamma$ to 77 decimal places (error $1.6\times 10^{-7}$), while ten terms of the original series (7) give only 11 correct digit. Twenty terms of (34) achieve $13$-digit accuracy. A numerical comparison appears in Section 8.
7.2 Tail bounds for direct computation of $h(k)$
For computational purposes, we bound the error incurred by truncating the defining series for $h(k)$.
Proposition 25.
Let $h_{N}(k):=\sum_{n=2}^{N}\operatorname{arctanh}(n^{-k})$. Then for $k\geq 2$ and $N\geq 2$,
$0<h(k)-h_{N}(k)\leq\frac{1}{1-(N+1)^{-2k}}\sum_{n=N+1}^{\infty}n^{-k}\leq\frac{1}{1-(N+1)^{-2k}}\!\left[(N\!+\!1)^{-k}+\frac{(N\!+\!1)^{1-k}}{k-1}\right].$ (36)
Proof.
Since $\operatorname{arctanh}(x)\leq x/(1-x^{2})$ for $0<x<1$ and $n^{-2k}\leq(N+1)^{-2k}$ for $n\geq N+1$,
$h(k)-h_{N}(k)$ $=\sum_{n=N+1}^{\infty}\operatorname{arctanh}(n^{-k})\leq\sum_{n=N+1}^{\infty}\frac{n^{-k}}{1-n^{-2k}}$
$\leq\frac{1}{1-(N+1)^{-2k}}\sum_{n=N+1}^{\infty}n^{-k}.$
The final inequality follows from the integral comparison $\sum_{n=N+1}^{\infty}n^{-k}\leq(N+1)^{-k}+\int_{N+1}^{\infty}x^{-k}\,dx$. ∎
Example 26.
For $k=2$ and $N=100$: the tail bound gives $h(2)-h_{100}(2)<1.0002\cdot(101^{-2}+101^{-1})<0.0100$. For N=10, [^0] N=10{,}000: the bound gives $h(2)-h_{10000}(2)<1.0001\cdot 10^{-4}$. In practice, the zeta-series representation (20) is preferable for high-precision computation, as it converges geometrically.
8 Numerical examples
We present numerical data illustrating the quality of our formulas.
8.1 Explicit values of $h(k)$
Using Theorem 1 and the product formulas from Theorem 5, we obtain exact closed-form expressions for all $k$.
$k$ $h(k)$ Formula type
2 $0.650923\ldots$ Elementary (hyperbolic)
3 $0.202733\ldots$ Gamma functions
4 $0.082405\ldots$ Trigonometric product
5 $0.036938\ldots$ Gamma functions
6 $0.017344\ldots$ Trigonometric product
8.2 Quality of the zeta approximation
$k$ $1+h(k)$ $C(k)$ Approx. Exact $\zeta(k)$
2 1.6509 0.0060 1.6449 1.6449
3 1.2027 0.0007 1.2021 1.2021
4 1.0824 0.0001 1.0823 1.0823
5 1.0369 0.0000 1.0369 1.0369
8.3 Convergence of the gamma representations
Table 1 compares the original representation (7) with the accelerated formula (34). The accelerated series achieves $13$-digit accuracy with only 2020 terms, while 100,000100{,}000 terms of the original series yield only 55 correct digits.
Terms NN Direct (7) Accelerated (34) Tail bound (35)
5 $0.587\ldots$ $0.577215664\ldots$ $2.03\times 10^{-5}$
10 $0.581\ldots$ $0.5772156649015\ldots$ $1.99\times 10^{-8}$
20 $0.579\ldots$ $0.5772156649015329\ldots$ $1.89\times 10^{-14}$
1,0001{,}000 $0.57894\ldots$ $0.5772156649015329\ldots$ $1.81454\times 10^{-604}$
10,00010{,}000 $0.57808\ldots$ $0.5772156649015329\ldots$ $5.23\times 10^{-6023}$
100,000100{,}000 $0.57722\ldots$ $0.5772156649015329\ldots$ $2.09\times 10^{-60208}$
Table 1: Comparison of convergence rates for the two gamma representations. The exact value is $\gamma=0.5772156649015329\ldots$ (A001620).
8.4 Verification of bounds
Table 2 illustrates the two-sided bounds from Theorem 19 for small values of $k$.
$k$ Lower: $\zeta(k)-1$ Exact $h(k)$ Upper bound (29)
2 $0.644934$ $0.650923$ $0.651317$
3 $0.202057$ $0.202733$ $0.202739$
4 $0.082323$ $0.082405$ $0.082406$
5 $0.036928$ $0.036938$ $0.036938$
Table 2: Two-sided bounds for $h(k)$. The gap between the lower and upper bounds narrows rapidly as $k$ increases.
8.5 OEIS sequence data
The decimal expansion of $h(2)=\frac{1}{2}\ln(\sinh(\pi)/\pi)$ appears in the OEIS as A372862. The Euler–Mascheroni constant $\gamma$ is A001620. The values of $\zeta(k)$ for small $k$ correspond to well-known constants: $\zeta(2)=\pi^{2}/6$ (A013661), $\zeta(3)$ is Apéry’s constant (A002117), $\zeta(4)=\pi^{4}/90$ (A013662), and so on.
The authors suggest that the decimal expansions of $h(k)$ for $k\geq 3$ and the correction values $C(k)$ may merit submission to the OEIS as new sequences.