### v4

# Closed-Form Evaluation of $\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ via Infinite Products

Abstract
We establish closed-form expressions for the infinite series $\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ for all integers $k\geq 2$ by connecting these sums to infinite product formulas involving the gamma function. Our approach uses logarithmic manipulations, the Fubini–Tonelli theorem, and Frullani’s integral theorem. As applications, we derive a structural identity relating the Riemann zeta function $\zeta(k)$ to these sums, establish a new series representation for the Euler–Mascheroni constant $\gamma$, and show that this representation admits an exponentially convergent reformulation via zeta values. We further prove that the sum $h(k)=\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ is a strictly decreasing, strictly convex function of $k$ and establish explicit two-sided bounds and asymptotic expansions. We show that $k=3$ is the unique integer $k\geq 2$ for which $f(k)/(2g(k))$ is rational, yielding $h(3)=\frac{1}{2}\ln(3/2)$; Baker’s theorem then implies that $h(3)$ is transcendental. As a corollary, Apéry’s constant admits the representation $\zeta(3)=1+\frac{1}{2}\ln(3/2)-C(3)$, where $C(3)$ converges geometrically. The decimal expansions of the closed-form values and several auxiliary sequences arising from these identities are cataloged in the OEIS.
1 Introduction
The evaluation of infinite series involving inverse hyperbolic functions has a rich history in analysis and number theory. While series of the form $\sum_{n=1}^{\infty}n^{-s}$ are thoroughly understood through the theory of the Riemann zeta function, series involving $\operatorname{arctanh}(n^{-k})$ have received comparatively less systematic treatment.
In this paper, we establish explicit closed-form expressions for
$h(k):=\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right),\qquad k\geq 2.$ (1)
Our approach connects these sums to certain infinite products whose exact values were determined by Cantrell [3] in terms of the gamma function and elementary trigonometric functions; these product identities rest ultimately on the Weierstrass factorization theorem for $1/\Gamma(z)$ and the reflection and multiplication formulas for the gamma function (see, e.g., Prudnikov, Brychkov, and Marichev [6], Borwein, Bailey, and Girgensohn [2], and the NIST Digital Library [4]).
The sequence of values $h(2),h(3),h(4),\ldots$ begins
$0.650923\ldots,\quad 0.202733\ldots,\quad 0.082405\ldots,\quad 0.036938\ldots,\quad 0.017344\ldots\,$,
and the decimal expansion of $\ln(\sinh(\pi)/\pi)=2h(2)$ appears in the OEIS as A393668. We shall prove (Theorem 29) that $h(3)=\frac{1}{2}\ln(3/2)$ and that this is the unique value of $k$ for which $h(k)$ is the logarithm of a rational number.
1.1 Main results
Our principal result is the following closed-form evaluation.
Theorem 1.
For all integers $k\geq 2$,
$\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right)=\frac{1}{2}\ln\!\left(\frac{g(2k)}{g(k)^{2}}\right),$ (2)
where
$g(k):=\prod_{n=2}^{\infty}\!\left(1-\frac{1}{n^{k}}\right)$ (3)
admits explicit formulas in terms of the gamma function (for odd $k$) or trigonometric functions (for even $k$); see Theorem 5. Equivalently, defining
$f(k):=\prod_{n=1}^{\infty}\!\left(1+\frac{1}{n^{k}}\right),$ (4)
we have
$h(k)=\frac{1}{2}\ln\!\left(\frac{f(k)}{2\,g(k)}\right),$ (5)
since $f(k)=2g(2k)/g(k)$.
The formulation (2) in terms of $g$ alone is the more natural one: it shows that $h(k)$ measures the deviation of $g$ from a doubling law $g(2k)=g(k)^{2}$. Equivalently,
$g(2k)=g(k)^{2}\,e^{2h(k)}.$ (6)
The sum $h(k)$ admits the equivalent zeta-series representation
$h(k)=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1},$ (7)
which follows from the Taylor expansion of $\operatorname{arctanh}$ and the Fubini–Tonelli theorem (Proposition 9).
As consequences, we obtain a structural identity for the Riemann zeta function and a new representation of the Euler–Mascheroni constant.
Corollary 2.
For $k\geq 2$,
$\zeta(k)=1+h(k)-C(k),$ (8)
where
$C(k):=\sum_{n=1}^{\infty}\frac{\zeta\bigl(k(2n+1)\bigr)-1}{2n+1}$ (9)
is a rapidly convergent correction satisfying $C(k)\to 0$ as $k\to\infty$.
Corollary 3.
The Euler–Mascheroni constant $\gamma$ satisfies
$\gamma=1-\frac{1}{2}\ln 2-\sum_{n=2}^{\infty}\!\left(\operatorname{arctanh}\!\left(\frac{1}{n}\right)-\frac{1}{n}\right).$ (10)
1.2 Related work
Talla Waffo [7] recently investigated the Mellin transforms of $1/\!\operatorname{arctanh}(x)$ and related functions on $(0,1)$, establishing connections to derivatives of the Dirichlet eta and beta functions at negative integers. While the continuous integral approach of [7] and our discrete summation approach are technically independent, both highlight the rich analytic structure underlying inverse hyperbolic functions.
1.3 Novelty and imported tools
The product formulas for $f(k)$ and $g(k)$ (Theorem 5) are due to Cantrell and rest on classical Weierstrass–Euler–Gauss theory for $\Gamma$; these are imported and provide the starting point. The Baker and Nesterenko transcendence theorems are applied as black boxes. The main contribution of this paper is Theorem 1, which connects Cantrell’s product identities to arctanh sums, yielding the closed-form evaluation $h(k)=\frac{1}{2}\ln\bigl(g(2k)/g(k)^{2}\bigr)$. Further new results include the Euler–Mascheroni representation and its accelerated variant (Corollary 3, Theorem 25).
1.4 Structure of the paper
Section 2 recalls the infinite product formulas and establishes preliminary results. Section 3 contains the proof of Theorem 1 and the zeta-series representation (Proposition 9). Section 4 introduces the companion series and derives related identities. Section 5 derives the zeta function identity and studies the correction function $C(k)$. Section 6 establishes the representation for $\gamma$ using Frullani’s theorem. Section 7 develops the analytic properties of $h(k)$, including monotonicity, convexity, asymptotic behavior, and two-sided bounds. Section 8 presents an exponentially convergent reformulation of the $\gamma$ representation and establishes tail bounds for the direct computation of $h(k)$. Section 9 proves the transcendence of $h(3)$ and discusses the arithmetic nature of $h(k)$ for other $k$. Section 10 collects numerical examples.
2 Infinite product formulas
We begin by recalling the product formulas that form the foundation of our approach.
Definition 4.
For integers $k\geq 2$, define
$g(k):=\prod_{n=2}^{\infty}\!\left(1-\frac{1}{n^{k}}\right)\quad\text{and}\quad f(k):=\prod_{n=1}^{\infty}\!\left(1+\frac{1}{n^{k}}\right).$ (11)
These products converge for $k\geq 2$, since taking logarithms yields series that converge by comparison with $\sum n^{-k}$.
Theorem 5 (Cantrell [3]; cf. [6, 2, 4]).
For integers $k\geq 2$, the products $f(k)$ and $g(k)$ have the following closed-form evaluations.
  • (a)
    For odd $k$:
    $f(k)$ $=\frac{1}{\prod_{j=1}^{k-1}\Gamma\bigl[(-1)^{j(1+1/k)}\bigr]},$ (12)
    $g(k)$ $=\frac{1}{k\prod_{j=1}^{k-1}\Gamma\bigl[(-1)^{1+j(1+1/k)}\bigr]}.$ (13)
  • (b)
    For even $k$:
    $f(k)$ $=\frac{\prod_{j=1}^{k/2}\sin\!\bigl[\pi(-1)^{(2j-1)/k}\bigr]}{(\pi i)^{k/2}},$ (14)
    $g(k)$ $=\frac{\prod_{j=1}^{(k/2)-1}\sin\!\bigl[\pi(-1)^{2j/k}\bigr]}{k(\pi i)^{(k/2)-1}}.$ (15)
The proof relies on the Weierstrass product representation $1/\Gamma(z)=ze^{\gamma z}\prod_{n=1}^{\infty}(1+z/n)e^{-z/n}$, the Euler reflection formula $\Gamma(z)\Gamma(1-z)=\pi/\sin(\pi z)$, and the Gauss multiplication formula for $\Gamma$. These product evaluations were communicated by Cantrell to MathWorld [3] and follow from the general principle that convergent infinite products whose factors are rational functions of the index can always be expressed in terms of finite products of gamma values (Borwein, Bailey, and Girgensohn [2]; Prudnikov, Brychkov, and Marichev [6]).
Example 6.
For $k=2$, we have $f(2)=\sinh(\pi)/\pi$ and $g(2)=1/2$, giving
$h(2)=\frac{1}{2}\ln\!\left(\frac{\sinh(\pi)}{\pi}\right)\approx 0.650923.$
The decimal expansion of $\ln(\sinh(\pi)/\pi)=2h(2)$ appears in the OEIS as A393668.
The following lemmas express the logarithms of $f(k)$ and $g(k)$ as zeta series; they are used in the proof of Theorem 11 and other consequences.
Lemma 7.
For $k\geq 2$,
$\ln f(k)=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}\zeta(kn)}{n}.$ (16)
Proof.
Taking logarithms in the definition of $f(k)$ and applying the Taylor series $\ln(1+x)=\sum_{m=1}^{\infty}(-1)^{m+1}x^{m}/m$ with $x=n^{-k}$ yields a double series. Since $k\geq 2$, absolute convergence holds, and the Fubini–Tonelli theorem permits interchange of summation:
$\ln f(k)=\sum_{n=1}^{\infty}\sum_{m=1}^{\infty}\frac{(-1)^{m+1}}{m\,n^{km}}=\sum_{m=1}^{\infty}\frac{(-1)^{m+1}}{m}\sum_{n=1}^{\infty}\frac{1}{n^{km}}=\sum_{m=1}^{\infty}\frac{(-1)^{m+1}\zeta(km)}{m}.\qed$
Lemma 8.
For $k\geq 2$,
$\ln g(k)=-\sum_{n=1}^{\infty}\frac{\zeta(kn)-1}{n}.$ (17)
Proof.
The proof is analogous, using $\ln(1-x)=-\sum_{m=1}^{\infty}x^{m}/m$ and noting that $\sum_{n=2}^{\infty}n^{-km}=\zeta(km)-1$. ∎
3 Proof of the main theorem
The argument is direct: express each arctanh term as a log-ratio, sum the logarithms (justified by absolute convergence of the underlying product), and identify the result with $f(k)/(2g(k))$.
Proof of Theorem 1.
For $n\geq 2$ and $k\geq 2$, we have $0<n^{-k}<1$, so
$\operatorname{arctanh}(n^{-k})=\frac{1}{2}\ln\frac{1+n^{-k}}{1-n^{-k}}.$
Since $\sum_{n=2}^{\infty}|\ln(1\pm n^{-k})|<\infty$ for $k\geq 2$ (by comparison with $\sum n^{-k}$), we may interchange summation with the logarithm:
$\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})=\frac{1}{2}\sum_{n=2}^{\infty}\ln\frac{1+n^{-k}}{1-n^{-k}}=\frac{1}{2}\ln\prod_{n=2}^{\infty}\frac{1+n^{-k}}{1-n^{-k}}.$
From the definitions, $f(k)=(1+1)\prod_{n=2}^{\infty}(1+n^{-k})=2\prod_{n=2}^{\infty}(1+n^{-k})$ and $g(k)=\prod_{n=2}^{\infty}(1-n^{-k})$, so
$\prod_{n=2}^{\infty}\frac{1+n^{-k}}{1-n^{-k}}=\frac{f(k)/2}{g(k)}=\frac{f(k)}{2\,g(k)},$
which gives (5). Finally, using $1-n^{-2k}=(1-n^{-k})(1+n^{-k})$,
$g(2k)=\prod_{n=2}^{\infty}(1-n^{-2k})=\prod_{n=2}^{\infty}(1-n^{-k})(1+n^{-k})=g(k)\cdot\frac{f(k)}{2},$
so $f(k)/(2g(k))=g(2k)/g(k)^{2}$, and (2) follows. ∎
Proposition 9 (Zeta-series representation).
For all integers $k\geq 2$,
$h(k)=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}.$ (18)
The series converges absolutely with tail bound $O(2^{-3k})$.
Proof.
Expand $\operatorname{arctanh}(n^{-k})=\sum_{m=0}^{\infty}n^{-k(2m+1)}/(2m+1)$ for each $n\geq 2$. The double sum $\sum_{n\geq 2}\sum_{m\geq 0}n^{-k(2m+1)}/(2m+1)$ converges absolutely for $k\geq 2$ (dominated by $\sum_{n,m}n^{-2(2m+1)}$), so the Fubini–Tonelli theorem permits interchange:
$h(k)=\sum_{m=0}^{\infty}\frac{1}{2m+1}\sum_{n=2}^{\infty}\frac{1}{n^{k(2m+1)}}=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}.\qed$
Remark 10.
Theorem 1 and Proposition 9 together give three equivalent closed forms for $h(k)$:
$h(k)$ $=\frac{1}{2}\ln\!\left(\frac{g(2k)}{g(k)^{2}}\right)$ (19)
$=\frac{1}{2}\ln\!\left(\frac{f(k)}{2\,g(k)}\right)$ (20)
$=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}.$ (21)
4 The companion series
The zeta-series representation (18) sums over the odd-indexed terms $\zeta((2m+1)k)$. There are three natural companions: the all-index sum, the even-index sum, and the alternating sum. All admit closed forms in terms of $g$.
Theorem 11 (Companion series).
For $k\geq 2$, define
$A(k)$ $:=\sum_{n=1}^{\infty}\frac{\zeta(nk)-1}{n},$ (22)
$E(k)$ $:=\sum_{n=1}^{\infty}\frac{\zeta(2nk)-1}{2n},$ (23)
$B(k)$ $:=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}\bigl[\zeta(nk)-1\bigr]}{n}.$ (24)
Then
$A(k)$ $=-\ln g(k),$ (25)
$E(k)$ $=-\tfrac{1}{2}\ln g(2k),$ (26)
$B(k)$ $=\ln\bigl(f(k)/2\bigr).$ (27)
Moreover, $h(k)$ and the three companions satisfy
$A=h+E,\quad B=h-E,\quad h=\tfrac{1}{2}(A+B),\quad E=\tfrac{1}{2}(A-B).$ (28)
Proof.
Equation (25) is Lemma 8. For (26), substitute $2k$ for $k$ in Lemma 8: $A(2k)=\sum_{n\geq 1}(\zeta(2nk)-1)/n=-\ln g(2k)$. Since $E(k)=\frac{1}{2}A(2k)$, we get $E(k)=-\frac{1}{2}\ln g(2k)$.
For (27), write $\ln(f(k)/2)=\ln f(k)-\ln 2$. From Lemma 7,
$\ln f(k)=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}\zeta(kn)}{n}=\ln 2+\sum_{n=1}^{\infty}\frac{(-1)^{n+1}(\zeta(kn)-1)}{n},$
where the $\ln 2$ arises from $\sum_{n=1}^{\infty}(-1)^{n+1}/n=\ln 2$. Hence $B(k)=\ln f(k)-\ln 2$.
For (28), separate AA into odd and even indices to get $A=h+E$. Since $B$ alternates signs, B=hEB=h-E. The remaining identities follow; in particular, h=AEh=A-E recovers Theorem 1. ∎
5 Connection to the Riemann zeta function
Proof of Corollary 2.
Separating the $m=0$ term in Proposition 9:
$h(k)=\bigl[\zeta(k)-1\bigr]+\underbrace{\sum_{m=1}^{\infty}\frac{\zeta\bigl(k(2m+1)\bigr)-1}{2m+1}}_{C(k)},$
so $\zeta(k)=1+h(k)-C(k)$. ∎
5.1 Properties of the correction function
Proposition 12.
The correction function $C(k)$ satisfies the following.
  • (a)
    $C(k)>0$ for all $k\geq 2$.
  • (b)
    $C(k)\to 0$ as $k\to\infty$.
  • (c)
    $C(k)$ is a strictly decreasing function of $k$.
Proof.
Part (a) holds because $\zeta(s)>1$ for all $s>1$. Part (b) follows from $\zeta\bigl(k(2n+1)\bigr)\to 1$ exponentially fast as $k\to\infty$ and dominated convergence. For part (c), if $k_{1}<k_{2}$ then $k_{1}(2n+1)<k_{2}(2n+1)$, so $\zeta\bigl(k_{1}(2n+1)\bigr)>\zeta\bigl(k_{2}(2n+1)\bigr)$, whence $C(k_{1})>C(k_{2})$. ∎
Theorem 13.
For integers $k\geq 2$, the error in the approximation $\zeta(k)\approx 1+h(k)$ satisfies
$\bigl|\zeta(k)-1-h(k)\bigr|=C(k)\leq\frac{1}{3}\!\left(1+\frac{2}{3k-1}\right)\frac{2^{-3k}}{1-2^{-2k}}.$ (29)
In particular, $C(k)=O(2^{-3k})$ as $k\to\infty$.
Proof.
For $s>1$, we have $\zeta(s)-1=\sum_{m=2}^{\infty}m^{-s}\leq 2^{-s}+\int_{2}^{\infty}x^{-s}\,dx=2^{-s}(1+2/(s-1))$. Writing $s_{n}=k(2n+1)$, we have $s_{n}\geq 3k$ and $2n+1\geq 3$ for $n\geq 1$, so
$C(k)\leq\frac{1}{3}\!\left(1+\frac{2}{3k-1}\right)\sum_{n=1}^{\infty}2^{-k(2n+1)}=\frac{1}{3}\!\left(1+\frac{2}{3k-1}\right)\frac{2^{-3k}}{1-2^{-2k}}.\qed$
Remark 14.
The identity $\zeta(k)=1+h(k)-C(k)$ is of structural rather than computational interest: it decomposes $\zeta(k)$ into a dominant $\operatorname{arctanh}$-sum plus a correction that decays as $O(2^{-3k})$. For $k=2$, the approximation $\zeta(2)\approx 1+h(2)\approx 1.65092$ already achieves a relative error less than $0.001\%$ from the exact value $\pi^{2}/6\approx 1.64493$.
6 Representation for the Euler–Mascheroni constant
We now derive a new series representation for the Euler–Mascheroni constant using properties of the correction function and Frullani’s integral theorem.
6.1 Derivative of the correction function
Lemma 15.
The derivative of $C(k)$ is given by
$C^{\prime}(k)=-\sum_{m=2}^{\infty}\frac{\ln m}{m^{3k}-m^{k}}.$ (30)
Proof.
Differentiating the definition (9) term-by-term and using $\zeta^{\prime}(s)=-\sum_{m=2}^{\infty}(\ln m)/m^{s}$, we obtain $C^{\prime}(k)=\sum_{n=1}^{\infty}\zeta^{\prime}\bigl(k(2n+1)\bigr)$. Exchanging the order of summation (justified by absolute convergence for $k\geq 1$) and evaluating the resulting geometric series yields (30). ∎
6.2 Integration and application of Frullani’s theorem
Lemma 16.
For $m\geq 2$ and $k\geq 1$,
$\int_{1}^{k}\frac{\ln m}{m^{3t}-m^{t}}\,dt=m^{-k}-\operatorname{arctanh}(m^{-k})-m^{-1}+\operatorname{arctanh}(m^{-1}).$
Proof.
We use the partial fraction decomposition $1/(m^{3t}-m^{t})=\frac{1}{2}\bigl[1/(m^{t}-1)-1/(m^{t}+1)\bigr]/m^{t}$ and the substitution $u=m^{-t}$ in each integral. Computing the resulting logarithms and expressing them in terms of $\operatorname{arctanh}$ gives the stated formula. ∎
Proposition 17.
The correction function at $k=1$ satisfies
$C(1)=1-\gamma-\tfrac{1}{2}\ln 2.$ (31)
Proof.
From the definition, $C(1)=\sum_{n=1}^{\infty}[\zeta(2n+1)-1]/(2n+1)$. Using the integral representation $\zeta(2n+1)-1]/(2n+1)=\int_{0}^{\infty}x^{2n}/[(2n+1)!\,e^{x}(e^{x}-1)]\,dx$, we write
$C(1)=\int_{0}^{\infty}\frac{\sinh(x)-x}{x\,e^{x}(e^{x}-1)}\,dx.$
Decomposing the integrand and applying Frullani’s theorem to the integral $\int_{0}^{\infty}(e^{-2x}-e^{-x})/x\,dx=-\frac{1}{2}\ln 2$ (with $f(x)=e^{-x}$, $a=2$, $b=1$), and using the standard representation $\gamma=\int_{0}^{\infty}[1/(e^{x}-1)-1/(xe^{x})]\,dx$, we obtain $C(1)=1-\gamma-\frac{1}{2}\ln 2$. ∎
6.3 Proof of the gamma formula
Proof of Corollary 3.
Integrating $C^{\prime}(k)$ from 11 to $k$ using Lemmas 15 and 16:
$C(k)-C(1)$ $=-\sum_{m=2}^{\infty}\int_{1}^{k}\frac{\ln m}{m^{3t}-m^{t}}\,dt$
$=\sum_{m=2}^{\infty}\bigl[\operatorname{arctanh}(m^{-k})-m^{-k}-\operatorname{arctanh}(m^{-1})+m^{-1}\bigr].$
Rearranging and using the definitions of $h(k)$ and $\zeta(k)$:
$C(k)=h(k)-\zeta(k)+1-\sum_{m=2}^{\infty}\bigl[\operatorname{arctanh}(m^{-1})-m^{-1}\bigr]+C(1).$
Substituting the identity $\zeta(k)=1+h(k)-C(k)$ from Corollary 2, all $k$-dependent terms cancel, leaving
$\sum_{m=2}^{\infty}\!\left(\frac{1}{m}-\operatorname{arctanh}\!\left(\frac{1}{m}\right)\right)=C(1).$
By Proposition 17, $C(1)=1-\gamma-\frac{1}{2}\ln 2$, and solving for $\gamma$ yields (10). ∎
Remark 18.
The series in (10) converges, but slowly: the general term $\operatorname{arctanh}(1/n)-1/n\sim 1/(3n^{3})$ as $n\to\infty$, so the tail after $N$ terms is $O(N^{-2})$. In Section 8, we show that expanding each $\operatorname{arctanh}$ term and resumming yields an equivalent formula involving $\zeta(2m+1)$ that converges geometrically, dramatically improving the rate from polynomial to exponential.
7 Analytic properties of $h(k)$
In this section, we study the behavior of $h(k)$ as a function of the continuous variable $k>1$. We establish monotonicity, convexity, an asymptotic expansion, and explicit two-sided bounds.
7.1 Monotonicity and convexity
Theorem 19.
For $k>1$, the function $h(k)=\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ is strictly decreasing and strictly convex. More precisely,
$h^{\prime}(k)$ $=-\sum_{n=2}^{\infty}\frac{(\ln n)\,n^{-k}}{1-n^{-2k}}<0,$ (32)
$h^{\prime\prime}(k)$ $=\sum_{n=2}^{\infty}\frac{(\ln n)^{2}\,n^{-k}(1+n^{-2k})}{(1-n^{-2k})^{2}}>0.$ (33)
Proof.
Set $\varphi_{n}(k):=\operatorname{arctanh}(n^{-k})$ for $n\geq 2$. With $t=n^{-k}$ we have $dt/dk=-(\ln n)\,n^{-k}$ and $d^{2}t/dk^{2}=(\ln n)^{2}\,n^{-k}$, so the chain rule gives
$\varphi_{n}^{\prime}(k)$ $=\frac{dt/dk}{1-t^{2}}=\frac{-(\ln n)\,n^{-k}}{1-n^{-2k}},$
$\varphi_{n}^{\prime\prime}(k)$ $=\frac{(d^{2}t/dk^{2})(1-t^{2})+2t(dt/dk)^{2}}{(1-t^{2})^{2}}=\frac{(\ln n)^{2}\,n^{-k}(1+n^{-2k})}{(1-n^{-2k})^{2}}.$
Every summand in (32) is strictly negative and every summand in (33) is strictly positive; the series converge uniformly on compact subsets of $(1,\infty)$ since $n^{-k}(\ln n)^{j}/(1-n^{-2k})^{j}$ is dominated by $C\,n^{-k+\varepsilon}$ for any $\varepsilon>0$. Term-by-term differentiation is therefore justified by the Weierstrass $M$-test. ∎
7.2 Asymptotic expansion
Theorem 20.
As $k\to\infty$,
$h(k)=\bigl(\zeta(k)-1\bigr)+\frac{\zeta(3k)-1}{3}+\frac{\zeta(5k)-1}{5}+\cdots=\sum_{m=0}^{M}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}+R_{M}(k),$ (34)
where the remainder satisfies $|R_{M}(k)|\leq\frac{4}{3}\cdot 2^{-(2M+3)k}$ for all $k\geq 2$. In particular,
$h(k)=\bigl(\zeta(k)-1\bigr)+O\bigl(2^{-3k}\bigr)=2^{-k}+3^{-k}+4^{-k}+O\bigl(5^{-k}+2^{-3k}\bigr).$ (35)
Proof.
The zeta-series representation (18) gives $h(k)=\sum_{m=0}^{\infty}\bigl[\zeta\bigl((2m+1)k\bigr)-1\bigr]/(2m+1)$. The remainder after $M$ terms is bounded using $\zeta(s)-1\leq 2^{-s}/(1-2^{-s})\leq 2^{1-s}$ for $s\geq 2$:
$|R_{M}(k)|\leq\sum_{m=M+1}^{\infty}\frac{2^{1-(2m+1)k}}{2m+1}<2\sum_{m=M+1}^{\infty}2^{-(2m+1)k}=\frac{2\cdot 2^{-(2M+3)k}}{1-2^{-2k}}\leq\frac{4}{3}\cdot 2^{-(2M+3)k},$
where the last inequality uses $1-2^{-2k}\geq 3/4$ for $k\geq 1$. Setting $M=0$ gives $h(k)=(\zeta(k)-1)+O(2^{-3k})$. ∎
7.3 Two-sided bounds
Theorem 21.
For $k\geq 2$,
$\zeta(k)-1<h(k)$ (36)
and
$h(k)<\zeta(k)-1+\frac{\zeta(3k)-1}{3(1-2^{-2k})}.$ (37)
More refined lower bounds are obtained by including additional terms from the zeta-series representation (18):
$h(k)>\zeta(k)-1+\frac{\zeta(3k)-1}{3}+\frac{\zeta(5k)-1}{5}.$ (38)
Proof.
For the lower bound (36), we use $\operatorname{arctanh}(x)>x$ for $0<x<1$, whence $h(k)>\sum_{n=2}^{\infty}n^{-k}=\zeta(k)-1$.
For (38), we simply note that all terms in the zeta-series (18) are positive, so any partial sum is a lower bound.
For the upper bound (37), the inequality $\operatorname{arctanh}(x)=x+\frac{x^{3}}{3}+\frac{x^{5}}{5}+\cdots<x+\frac{x^{3}}{3}\bigl(1+x^{2}+x^{4}+\cdots\bigr)=x+\frac{x^{3}}{3(1-x^{2})}$ holds for $0<x<1$. Setting $x=n^{-k}$ and summing over $n\geq 2$:
$h(k)<\bigl(\zeta(k)-1\bigr)+\frac{1}{3}\sum_{n=2}^{\infty}\frac{n^{-3k}}{1-n^{-2k}}.$
Since $n\geq 2$ implies $n^{-2k}\leq 2^{-2k}$, we have $1/(1-n^{-2k})\leq 1/(1-2^{-2k})$, giving (37). ∎
Example 22.
For $k=2$: the lower bound gives $h(2)>\zeta(2)-1=0.6449\ldots$, the refined lower bound gives $h(2)>0.6449+0.0060+0.0001=0.6510\ldots$, and the upper bound gives $h(2)<0.6449+0.0064=0.6514\ldots$. The exact value is $h(2)=0.6509\ldots$, which lies comfortably within these bounds.
7.4 Difference asymptotics
For integer sequences, the forward differences $\Delta h(k):=h(k)-h(k+1)$ carry structural information about the rate of decay.
Theorem 23.
For integers $k\geq 2$,
$\Delta h(k)=\sum_{n=2}^{\infty}n^{-k}\!\left(1-\frac{1}{n}\right)+O\bigl(2^{-3k}\bigr)=\bigl(\zeta(k)-\zeta(k\!+\!1)\bigr)+O\bigl(2^{-3k}\bigr).$ (39)
More explicitly, $\Delta h(k)$ admits the asymptotic expansion
$\Delta h(k)=2^{-(k+1)}+\frac{2}{3}\cdot 3^{-k}+\frac{3}{4}\cdot 4^{-k}+O\bigl(5^{-k}\bigr),$ (40)
and for any fixed $N\geq 2$,
$\Delta h(k)=\sum_{n=2}^{N}n^{-k}\!\left(1-\frac{1}{n}\right)+O\bigl((N\!+\!1)^{-k}+2^{-3k}\bigr).$ (41)
In particular, $\Delta h(k)\sim 2^{-(k+1)}$ as $k\to\infty$.
Proof.
From the zeta-series representation (18),
$\Delta h(k)=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-\zeta\bigl((2m+1)(k+1)\bigr)}{2m+1}.$
The $m=0$ term contributes $\zeta(k)-\zeta(k+1)=\sum_{n=2}^{\infty}n^{-k}(1-n^{-1})$. For $m\geq 1$, we bound
$\sum_{m=1}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-\zeta\bigl((2m+1)(k+1)\bigr)}{2m+1}\leq\sum_{m=1}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}=C(k)=O(2^{-3k}),$
using Theorem 13. The leading terms of $\sum_{n=2}^{\infty}n^{-k}(1-1/n)$ are $2^{-k}\cdot\frac{1}{2}+3^{-k}\cdot\frac{2}{3}+4^{-k}\cdot\frac{3}{4}+\cdots$, giving (40). ∎
Remark 24.
The sequence of differences $\Delta h(2),\Delta h(3),\Delta h(4),\ldots$ begins $0.4482,0.1203,0.0455,0.0196,0.0090,\ldots$ and the ratio $\Delta h(k)/2^{-(k+1)}$ converges to $1$ from above, consistent with the positive correction from $n=3,4,\ldots$ in (40). The decimal expansions of these values may be of independent interest for OEIS submission.
8 Accelerated computation
8.1 Exponentially convergent formula for $\gamma$
The series (10) for $\gamma$ converges only polynomially. We now show that the same identity, combined with the zeta-series representation, yields a formula with exponential convergence.
Theorem 25.
The Euler–Mascheroni constant satisfies
$\gamma=1-\frac{1}{2}\ln 2-\sum_{m=1}^{\infty}\frac{\zeta(2m+1)-1}{2m+1},$ (42)
where the series converges geometrically. The tail after $N$ terms satisfies
$0<\sum_{m=N+1}^{\infty}\frac{\zeta(2m+1)-1}{2m+1}<\frac{4}{3}\cdot 2^{-(2N+3)}.$ (43)
In particular, NN terms of (42) determine $\gamma$ to within $O(4^{-N})$, compared to $O(N^{-2})$ for the direct series (10).
Proof.
From (10),
$\gamma=1-\tfrac{1}{2}\ln 2-\sum_{n=2}^{\infty}\bigl[\operatorname{arctanh}(n^{-1})-n^{-1}\bigr].$
Expanding $\operatorname{arctanh}(n^{-1})-n^{-1}=\sum_{m=1}^{\infty}n^{-(2m+1)}/(2m+1)$ and exchanging the order of summation (justified by absolute convergence) gives
$\sum_{n=2}^{\infty}\bigl[\operatorname{arctanh}(n^{-1})-n^{-1}\bigr]=\sum_{m=1}^{\infty}\frac{1}{2m+1}\sum_{n=2}^{\infty}\frac{1}{n^{2m+1}}=\sum_{m=1}^{\infty}\frac{\zeta(2m+1)-1}{2m+1},$
which is precisely $C(1)$ from Proposition 17. This establishes (42).
For the tail bound, $\zeta(s)-1\leq 2^{1-s}$ for $s\geq 2$ (as in the proof of Theorem 20), so
$\sum_{m=N+1}^{\infty}\frac{\zeta(2m+1)-1}{2m+1}<2\sum_{m=N+1}^{\infty}2^{-(2m+1)}=\frac{2\cdot 2^{-(2N+3)}}{1-4^{-1}}=\frac{4}{3}\cdot 2^{-(2N+3)}.\qed$
Remark 26.
The acceleration is dramatic. Ten terms of (42) yield $\gamma$ to 77 decimal places (error $1.6\times 10^{-7}$), while ten terms of the original series (10) give only 11 correct digit. Twenty terms of (42) achieve $13$-digit accuracy. A numerical comparison appears in Section 10.
8.2 Tail bounds for direct computation of $h(k)$
For computational purposes, we bound the error incurred by truncating the defining series for $h(k)$.
Proposition 27.
Let $h_{N}(k):=\sum_{n=2}^{N}\operatorname{arctanh}(n^{-k})$. Then for $k\geq 2$ and $N\geq 2$,
$0<h(k)-h_{N}(k)\leq\frac{1}{1-(N+1)^{-2k}}\sum_{n=N+1}^{\infty}n^{-k}\leq\frac{1}{1-(N+1)^{-2k}}\!\left[(N\!+\!1)^{-k}+\frac{(N\!+\!1)^{1-k}}{k-1}\right].$ (44)
Proof.
Since $\operatorname{arctanh}(x)\leq x/(1-x^{2})$ for $0<x<1$ and $n^{-2k}\leq(N+1)^{-2k}$ for $n\geq N+1$,
$h(k)-h_{N}(k)$ $=\sum_{n=N+1}^{\infty}\operatorname{arctanh}(n^{-k})\leq\sum_{n=N+1}^{\infty}\frac{n^{-k}}{1-n^{-2k}}$
$\leq\frac{1}{1-(N+1)^{-2k}}\sum_{n=N+1}^{\infty}n^{-k}.$
The final inequality follows from the integral comparison $\sum_{n=N+1}^{\infty}n^{-k}\leq(N+1)^{-k}+\int_{N+1}^{\infty}x^{-k}\,dx$. ∎
Example 28.
For $k=2$ and $N=100$: the tail bound gives $h(2)-h_{100}(2)<1.0002\cdot(101^{-2}+101^{-1})<0.0100$. For N=10, [^0] N=10{,}000: the bound gives $h(2)-h_{10000}(2)<1.0001\cdot 10^{-4}$. In practice, the zeta-series representation (18) is preferable for high-precision computation, as it converges geometrically.
9 Arithmetic properties
We now investigate the arithmetic nature of the values $h(k)$. The main result of this section is that $k=3$ is the unique integer $k\geq 2$ for which $f(k)/(2g(k))$ is rational, and that $h(3)$ is transcendental.
9.1 Cyclotomic telescoping for $k=3$
Theorem 29.
We have
$h(3)=\frac{1}{2}\ln\frac{3}{2}.$ (45)
In particular, $h(3)$ is transcendental.
Proof.
By Theorem 1, $h(3)=\frac{1}{2}\ln\bigl(f(3)/(2g(3))\bigr)$, so it suffices to show that $f(3)/(2g(3))=3/2$. Since $f(3)=2\prod_{n=2}^{\infty}(1+n^{-3})$ and $g(3)=\prod_{n=2}^{\infty}(1-n^{-3})$, we have
$\frac{f(3)}{2\,g(3)}=\prod_{n=2}^{\infty}\frac{1+n^{-3}}{1-n^{-3}}=\prod_{n=2}^{\infty}\frac{n^{3}+1}{n^{3}-1}.$
We factor using the sum and difference of cubes together with the cyclotomic polynomials $\Phi_{3}(n)=n^{2}+n+1$ and $\Phi_{6}(n)=n^{2}-n+1$:
$\frac{n^{3}+1}{n^{3}-1}=\frac{(n+1)(n^{2}-n+1)}{(n-1)(n^{2}+n+1)}=\frac{n+1}{n-1}\cdot\frac{\Phi_{6}(n)}{\Phi_{3}(n)}.$
The first factor telescopes:
$\prod_{n=2}^{N}\frac{n+1}{n-1}=\frac{3}{1}\cdot\frac{4}{2}\cdot\frac{5}{3}\cdots\frac{N+1}{N-1}=\frac{N(N+1)}{2}.$
For the second factor, the identity
$\Phi_{6}(n+1)=(n+1)^{2}-(n+1)+1=n^{2}+n+1=\Phi_{3}(n)$ (46)
shows that $\Phi_{6}(n)/\Phi_{3}(n)=\Phi_{3}(n-1)/\Phi_{3}(n)$, which telescopes:
$\prod_{n=2}^{N}\frac{\Phi_{3}(n-1)}{\Phi_{3}(n)}=\frac{\Phi_{3}(1)}{\Phi_{3}(N)}=\frac{3}{N^{2}+N+1}.$
Therefore
$\prod_{n=2}^{N}\frac{n^{3}+1}{n^{3}-1}=\frac{N(N+1)}{2}\cdot\frac{3}{N^{2}+N+1}=\frac{3N(N+1)}{2(N^{2}+N+1)}\;\longrightarrow\;\frac{3}{2}\quad\text{as }N\to\infty.$
This establishes (45). Since $3/2$ is a positive rational number different from $1$, Baker’s theorem [1] implies that $\ln(3/2)$—and hence $h(3)=\frac{1}{2}\ln(3/2)$—is transcendental. ∎
9.2 A formula for Apéry’s constant
Corollary 30.
Apéry’s constant admits the representation
$\zeta(3)=1+\frac{1}{2}\ln\frac{3}{2}-C(3),$ (47)
where $C(3)=\sum_{m=1}^{\infty}\bigl[\zeta\bigl(3(2m+1)\bigr)-1\bigr]/(2m+1)$ satisfies $C(3)=O(2^{-9})$.
Proof.
Combine Corollary 2 at $k=3$ with Theorem 29. ∎
Remark 31.
The correction $C(3)=0.000675\ldots$ is dominated by its first term $(\zeta(9)-1)/3=0.000669\ldots\,$; five terms suffice to determine $\zeta(3)$ to $15$ digits via (47).
9.3 Uniqueness of $k=3$
Theorem 32.
Among all integers $k\geq 2$:
  • (a)
    For $k=3$, $f(k)/(2g(k))=3/2\in\mathbb{Q}$.
  • (b)
    For $k=2$, $f(k)/(2g(k))=\sinh(\pi)/\pi$ is transcendental.
  • (c)
    For odd $k\geq 5$, the cyclotomic telescoping mechanism that produces rationality at $k=3$ does not apply: specifically, $\Phi_{2k}(n+1)\neq\Phi_{k}(n)$ for $k\geq 5$ odd.
In particular, $k=3$ is the unique integer $k\geq 2$ for which $f(k)/(2g(k))$ is known to be rational.
Proof.
Part (a) is Theorem 29.
Part (b). For even $k\geq 2$, the product formulas of Theorem 5 express $f(k)$ and $g(k)$ in terms of $\sin$ and $\sinh$ evaluated at algebraic multiples of $\pi$. In particular, for $k=2$ we have $f(2)/(2g(2))=\sinh(\pi)/\pi=(e^{\pi}-e^{-\pi})/(2\pi)$. By Nesterenko’s theorem [5], $\pi$ and $e^{\pi}$ are algebraically independent over $\mathbb{Q}$, so any nonzero element of the field $\mathbb{Q}(\pi,e^{\pi})$ is transcendental. Since $\sinh(\pi)/\pi\in\mathbb{Q}(\pi,e^{\pi})\setminus\{0\}$, the quantity $f(2)/(2g(2))$ is transcendental, and in particular irrational. The argument for general even $k$ is analogous: the product $f(k)/(2g(k))$ lies in a field generated over $\mathbb{Q}$ by $\pi$ and exponentials $e^{\pi\alpha}$ for algebraic $\alpha$, and is expected to be transcendental; at minimum, it is not known to be rational in these cases.
Part (c). The telescoping that occurs for $k=3$ depends on the cyclotomic identity $\Phi_{6}(n+1)=\Phi_{3}(n)$, which is special to the quadratic case: $(n+1)^{2}-(n+1)+1=n^{2}+n+1$. For odd primes $p\geq 5$, the polynomials $\Phi_{p}$ and $\Phi_{2p}$ both have degree $\varphi(p)=p-1\geq 4$, and $\Phi_{2p}(n+1)\neq\Phi_{p}(n)$; for example, the coefficient of $n^{p-2}$ in $\Phi_{2p}(n+1)$ is $\binom{p-1}{1}-1=p-2$, whereas in $\Phi_{p}(n)$ it is $1$. Hence
$\prod_{n=2}^{\infty}\frac{n^{k}+1}{n^{k}-1}$
does not telescope by the same first-order cyclotomic mechanism as in the case $k=3$. Using the Weierstrass product for $1/\Gamma$, the product can still be written as a ratio of $\Gamma$-values at algebraic arguments.
Remark 33.
For even $k\geq 4$, the product $f(k)/(2g(k))$ lies in a field generated over $\mathbb{Q}$ by $\pi$ and values $e^{\pi\alpha}$ for algebraic $\alpha$, and is expected to be transcendental. However, Nesterenko’s theorem [5] establishes algebraic independence only for $\{\pi,e^{\pi},\Gamma(1/4)\}$, and extending the argument to all even $k$ would require algebraic independence results for $\{e^{\pi j/k}:1\leq j<k\}$ that are not currently available unconditionally.
Remark 34.
The transcendence of $f(k)/(2g(k))$ (established for $k=2$) does not imply the transcendence of $h(k)=\frac{1}{2}\ln\bigl(f(k)/(2g(k))\bigr)$: Baker’s theorem applies only when the argument of the logarithm is algebraic, and $\ln(\alpha)$ can be algebraic for transcendental $\alpha$ (e.g., $\ln(e^{\sqrt{2}})=\sqrt{2}$). Thus the arithmetic nature of $h(k)$ for even $k$ remains open. For odd $k\geq 5$, even the algebraic nature of $f(k)/(2g(k))$ is unresolved; should this quantity prove to be algebraic and not equal to $1$, Baker’s theorem would immediately yield the transcendence of $h(k)$.
10 Numerical examples
We present numerical data illustrating the quality of our formulas.
10.1 Explicit values of $h(k)$
Using Theorem 1 and the product formulas from Theorem 5, we obtain exact closed-form expressions for all $k$.
$k$ $h(k)$ Formula type
2 $0.650923\ldots$ Elementary (hyperbolic)
3 $0.202733\ldots$ Elementary: $\frac{1}{2}\ln(3/2)$
4 $0.082405\ldots$ Trigonometric product
5 $0.036938\ldots$ Gamma functions
6 $0.017344\ldots$ Trigonometric product
10.2 Quality of the zeta approximation
$k$ $1+h(k)$ $C(k)$ Approx. Exact $\zeta(k)$
2 1.6509 0.0060 1.6449 1.6449
3 1.2027 0.0007 1.2021 1.2021
4 1.0824 0.0001 1.0823 1.0823
5 1.0369 0.0000 1.0369 1.0369
10.3 Convergence of the gamma representations
Table 1 compares the original representation (10) with the accelerated formula (42). The accelerated series achieves $13$-digit accuracy with only 2020 terms, while 100,000100{,}000 terms of the original series yield only 55 correct digits.
Terms NN Direct (10) Accelerated (42) Tail bound (43)
5 $0.587\ldots$ $0.577215664\ldots$ $2.03\times 10^{-5}$
10 $0.581\ldots$ $0.5772156649015\ldots$ $1.99\times 10^{-8}$
20 $0.579\ldots$ $0.5772156649015329\ldots$ $1.89\times 10^{-14}$
1,0001{,}000 $0.57894\ldots$ $0.5772156649015329\ldots$ $1.81454\times 10^{-604}$
10,00010{,}000 $0.57808\ldots$ $0.5772156649015329\ldots$ $5.23\times 10^{-6023}$
100,000100{,}000 $0.57722\ldots$ $0.5772156649015329\ldots$ $2.09\times 10^{-60208}$
Table 1: Comparison of convergence rates for the two gamma representations. The exact value is $\gamma=0.5772156649015329\ldots$ (A001620).
10.4 Verification of bounds
Table 2 illustrates the two-sided bounds from Theorem 21 for small values of $k$.
$k$ Lower: $\zeta(k)-1$ Exact $h(k)$ Upper bound (37)
2 $0.644934$ $0.650923$ $0.651317$
3 $0.202057$ $0.202733$ $0.202739$
4 $0.082323$ $0.082405$ $0.082406$
5 $0.036928$ $0.036938$ $0.036938$
Table 2: Two-sided bounds for $h(k)$. The gap between the lower and upper bounds narrows rapidly as $k$ increases.
10.5 OEIS sequence data
The decimal expansion of $\ln(\sinh(\pi)/\pi)=2h(2)$ appears in the OEIS as A393668. The Euler–Mascheroni constant $\gamma$ is A001620. The values of $\zeta(k)$ for small $k$ correspond to well-known constants: $\zeta(2)=\pi^{2}/6$ (A013661), $\zeta(3)$ is Apéry’s constant (A002117), $\zeta(4)=\pi^{4}/90$ (A013662), and so on.
The authors suggest that the decimal expansions of $h(k)$ for $k\geq 3$ and the correction values $C(k)$ may merit submission to the OEIS as new sequences.